OpenStudy (anonymous):

Find the slope of the line described by 8x = 48 + 3y

OpenStudy (anonymous):

Slope Intercept y = mx + b where: m is the slope b is the y-intercept

OpenStudy (anonymous):

First, we solve for y. So here y=(-8/3)x-16. Since y=mx+b where m is the slope, we know the slope here to be (-8/3)

OpenStudy (anonymous):